OpenStudy (anonymous):

given that cos (A+B) = cosAcosB-sinAsinB and f(Q)=8sin^4(θ)-2sin^2(θ)-2 show that f(θ)=cos4θ-3cos2θ Now solve, 8sin^4(θ)-2sin^2θ+3cos2θ=2.5 for 0<θ

OpenStudy (anonymous):

oh geez...I'm not good at this but Sarkar is. Maybe you could msg him?

OpenStudy (anonymous):

just go to the envelope on the top left and put Sarkar in the recipients box :) and the link to this problem!

OpenStudy (anonymous):

Idk either...probably the person who created this site.

myininaya (myininaya):

@omar1131 I'm confused a little about your problem

myininaya (myininaya):

f(Q)=8sin^4(θ)-2sin^2(θ)-2 show that f(θ)=cos4θ-3cos2θ

myininaya (myininaya):

So Q or theta for the first function?

myininaya (myininaya):

\[8 \sin^4(\theta)-2 \sin^2(\theta)-2=\cos(4 \theta) -3 \cos(2 \theta)\] So we want to show this is an identity right?

myininaya (myininaya):

assuming it is anyways...

myininaya (myininaya):

so notice that on on side we have are angles being double and 2*double lets try to right without the double thing going on

myininaya (myininaya):

\[\cos(4 \theta) -3 \cos(2 \theta)= \cos(2 \theta+2 \theta)-3 \cos( \theta+\theta)\] \[=(\cos(2 \theta) \cos(2 \theta)-\sin(2 \theta) \sin(2 \theta))-3(\cos(\theta) \cos(\theta)-\sin(\theta) \sin(\theta))\] I used the thing you gave us okay ? :) \[=\cos( 2 \theta) \cos( 2 \theta) -\sin(2 \theta) \sin(2 \theta))-3 \cos^2(\theta)+3 \sin^2(\theta)\] ------------------------------------------------- Now we just found that \[\cos(2 \theta)=\cos(\theta+\theta)=\cos^2(\theta)-\sin^2(\theta)\] But we still have some double action going on with the \[\sin(2 \theta)\] Can we use anything else besides the thing given? ------------------------------------------------- \[(\cos^2(\theta)-\sin^2(\theta)(\cos^2(\theta)-\sin^2(\theta))-\sin(2\theta) \sin(2 \theta)-3 \cos^2(\theta)+3 \sin^2(\theta)\] \[=\cos^4(\theta)-2 \cos^2(\theta) \sin^2(\theta) +\sin^4(\theta)-\sin(2 \theta) \sin(2 \theta)-3 \cos^2(\theta)+3 \sin^2(\theta)\]

myininaya (myininaya):

So can we use anything else besides the thing given?

myininaya (myininaya):

Because I can think of another useful sum/difference formula

myininaya (myininaya):

like can we use \[\sin(2 \theta)=2 \sin(\theta) \cos(\theta)\]

OpenStudy (anonymous):

maybe use \[\cos(2\theta) = 1 - 2\sin^2(\theta)\]

OpenStudy (anonymous):

rearranging to get \[ -2\sin^2(\theta) \] and \[8\sin^4(\theta)\] and substitute those in

OpenStudy (anonymous):

\[\cos(2\theta) = \cos^2(\theta) - \sin^2(\theta) = 2\cos^2(\theta) -1 = 1 - 2\sin^2(\theta)\]

OpenStudy (anonymous):

substituting in those in my earlier comment leaves you with \[2(1 - \cos(2\theta))^2 + \cos(2\theta) -3\]

OpenStudy (anonymous):

try multiplying out the first bracket, collecting like terms and seeing if you can spot the other double angle identity lurking around somewhere...

OpenStudy (anonymous):

hope this is helping xD

OpenStudy (anonymous):

i can post up a detailed solution for reference if you want

myininaya (myininaya):

@omar1131 I understand your anger. But OpenStudy is a pg13 website. Please don't use the bad language here.

myininaya (myininaya):

I would still like to know the answer to my one question Can we use \[\sin(2 \theta)=2 \sin(\theta) \cos(\theta)\]

myininaya (myininaya):

Ok great! :)

myininaya (myininaya):

\[ =\cos^4(\theta)-2 \cos^2(\theta) \sin^2(\theta) +\sin^4(\theta)-\sin(2 \theta) \sin(2 \theta)-3 \cos^2(\theta)+3 \sin^2(\theta) \] \[=\cos^4(\theta)-2 \cos^2(\theta) \sin^2(\theta)+\sin^4(\theta)-2 \sin(\theta) \cos(\theta) 2 \sin(\theta) \cos(\theta)-3 \cos^2(\theta)+3 \sin^2(\theta)\] \[=\cos^4(\theta)-2 \cos^2(\theta) \sin^2(\theta)+\sin^4(\theta)-2 \sin^2(\theta)\cos^2(\theta)-3 \cos^2(\theta)+3 \sin^2(\theta)\] \[=\cos^4(\theta)-4 \sin^2(\theta) \cos^2(\theta)-3\cos^2(\theta) +3 \sin^2(\theta)+\sin^4(\theta)\] Now the other side is in terms of sine so lets write this side in terms of sine

OpenStudy (anonymous):

we have \[\cos(A + B) = cosAcosB - sinAsinB\] so\[\cos(2\theta) = \cos(\theta + \theta) = \cos(\theta)\cos(\theta) - \sin(\theta)\sin(\theta)\]\[= \cos^2(\theta) -\sin^2(\theta) [1]\] now using\[\cos^2(\theta) +\sin^2(\theta) = 1[2]\]\[\sin^2(\theta) = (1-\cos^2(\theta)) [3]\] substitute equation [3] into equation [1]\[\cos(2\theta) = \cos^2(\theta) - (1-\cos^2(\theta))\]\[\cos(2\theta)= 2\cos^2(\theta) -1\] doing the same but rearranging [2] to get cos^2 and subbing we get \[\cos(2\theta)= 1 - 2sin^2(\theta)\]

OpenStudy (anonymous):

ill let myininaya finish, thats a good way (although i prefer my solution :p)

OpenStudy (anonymous):

cya guys

myininaya (myininaya):

\[=(\cos^2(\theta))^2-4\sin^2(\theta)(1-\sin^2(\theta))-3(1-\sin^2(\theta))+3 \sin^2(\theta)+\sin^4(\theta)\] \[=(1-\sin^2(\theta))^2-4\sin^2(\theta)+4 \sin^4(\theta)-3 +3 \sin^2(\theta)+3 \sin^2(\theta)+\sin^4(\theta)\] \[=1-2\sin^2(\theta)+\sin^4(\theta)-4\sin^2(\theta)+4 \sin^4(\theta)-3 +3 \sin^2(\theta)+3 \sin^2(\theta)+\sin^4(\theta)\] Like terms :) \[=6 \sin^4(\theta) +0\sin^2(\theta)-2 =6 \sin^4(\theta)-2 \] I think I made a mistake somewhere :(

myininaya (myininaya):

I need to do this by hand lol

myininaya (myininaya):

lol ok @omar1131 but let me know if you run into any more trouble This one is crazy long

myininaya (myininaya):

Crazy long the way I'm approaching it lol

myininaya (myininaya):



myininaya (myininaya):

ok i got it using the the identities i was talking about

myininaya (myininaya):

I think I will just scan this and let you look at it is that okay?

myininaya (myininaya):

I guess you can say that lol But we are just suppose to show they are equal and that is what I did Or that is what I'm assume from asking earlier

myininaya (myininaya):

Well we showed that \[\cos(4 \theta)-3 \cos(2 \theta)=8 \sin^4(\theta)-2 \sin^2(\theta)-2\] But we know that \[f( \theta)=8 \sin^4(\theta)-2 \sin^2(\theta)-2 \] And since \[\cos(4 \theta)-3 \cos(2 \theta)=8 \sin^4(\theta)-2 \sin^2(\theta)-2\] then we also have \[f(\theta)=\cos(4 \theta)-3 \cos(2 \theta)\] And we are done with this part of your problem(s) lol well @omar1131 Some people may feel uncomfortable about give their private info out(e.x their name) so it is against the CoC. But my name is Christy :)

myininaya (myininaya):

It is nice to meet you to and I'm sorry you are confused. What part confuses you exactly?